A store purchased an item for $90 and planned to sell it for $153.00 so that their profit would be 50% of their cost. If they were unable to sell it for this amount, what minimum selling price would allow them to break even
Found 2 solutions by math_tutor2020, greenestamps:
Answer by math_tutor2020(3816)


I don't know where your teacher is getting the "profit would be 50% of their cost"

The profit itself is
profit = revenue - cost = 153 - 90 = 63
which leads to
profit/cost = 63/90 = 0.70 = 70%

So the "50%" should be "70%".

If the shop owner wants to break even, then they would sell the item at the price they purchased it.
Meaning the sale price would be $90.
The breakeven point is when the profit is 0.

Answer by greenestamps(13198)



Look closely at what you are posting before you post it. If the item was bought for $90 and the profit was 50%, then the selling price is $135, not $153.

Trivially, to break even the store must sell it for the same price they paid for it, which is $90.

Note that the planned profit and planned selling price are completely irrelevant to the question that is asked....